Solve and graph the solution set:
A. 5x-4(6+2x)<2(1-x)
B. x-9<3x-4 C. 12-/7-2m/=9
On C. the // means absolute value

Answer by RAY100(1637)

(A) 5x-4(6+2x)<2(1-x)
5x -24 -8x <2-2x
-1x < 26
x> -26
.
check x=-25,,5(-25)-4(6+2(-25)) < 2(1-(-25)
.
-125+176< 52
.
51<52,,,,,,ok
.
(B) x-9<3x-4
-5 < 2x
-5/2 < x
.
check,,,let x=-2,,,,(-2)-9<3(-2)-4
-11<-10,,,,ok
.
.(C) 12-[7-2m] =9
12-9=[7-2m]
3=[7-2m]
.
(7-2m) = +3,,,,,,,,,,,,,,,,,(7-2m)= -3
-2m = -4,,,,,,,,,,,,,,,,,,,,,,,-2m = -10
m=2,,,,,,,,,,,,,,,,,,,,,,,,,,,m=5
.
check,,(2) 12-[7-2(2)] =9,,,,,12-[+3]=9,,,12-3=9,,,,ok
.
(5),,,12-[7-2(5)]=9,,,,,12-[-3]=9,,,12-3=9,,,,ok